# Disentangling Exploration of Large Language Models by Optimal Exploitation

### v1

@@ 6 Conclusion
In summary, with this work, we proposed a decomposition based on the optimal exploitation return to evaluate LLM-based agents for intelligent state-space exploration. Our evaluation includes an adaption of a popular environment from the reinforcement learning domain that goes beyond stateless multi-armed bandits and is less domain biased than other popular benchmarks. We believe developing systems that can do both, intelligent exploitation and exploration, are a crucial step towards truly generalist agents. Future work could apply our framework to evaluate agents in continuous and more complex environments. For example, we plan to include more randomization, stochasticity, and task-specific exploration. A limitation of our domain is that rewards are randomly spread and high return, therefore, correlates with high state-space coverage. By developing environments that combine complex exploration with underlying rules, domain knowledge and different tasks, we could use our framework to test more intelligent exploration of agents. We believe, as exploitation and exploration are both difficult problems, it is essential to evaluate both in isolation.
Appendix A Environment
@@
A.1 Layouts
Figure 4: 4x4 grid variations.
Figure 5: 5x5 grid variations.
Figure 6: 7x7 grid variations.
A.2 Prompt Template
Large Language Model Agent Your past episodes: <HISTORY> You are controlling an agent in an unknown world. Over a total of 20 episodes, you can interact with objects in the environment. <INSTRUCTION> You have <MAX STEPS> door interactions per episode but can pick up three balls, keys, or boxes. Current episode: <EPISODIC HISTORY> Which object do you want to interact with next? Reply with one object enclosed with <and >, e.g. <door>. What is your next action?
Figure 7: Prompt Template
A.3 Abstraction Wrapper
@@
Our room worlds are built upon the minigrid framework. To address the limitations of the low-level action and observation spaces, we use an abstraction wrapper function. This function transforms raw observations and actions into higher-level abstractions, allowing agents to reason more effectively about interactions with objects and doors within a room. By abstracting low-level details, the evaluation focuses on the agent’s exploration strategy rather than basic navigation capabilities. The abstraction wrapper defines the state as a discrete set of available objects and doors within the current room. Each room is uniquely configured to ensure that the agent can differentiate between them. Agent actions are simplified to interactions with specific objects or doors (e.g., choosing to open a door or collect an object). Once the agent selects an action at this abstract level, the wrapper translates it into a sequence of low-level actions using a predefined low-level policy to execute the selected behavior. Our wrapper can be seen as bridging the minigrid framework to a representation similar to text-environments like TextWorld.
A.4 Optimal Exploitation
@@
We formulate the computation of the optimal exploitation return $R^{opt}$ as an instance of the Orienteering Problem. We represent the explored balls in the state-space of history $h_{t}$ as a graph $G=(V,E)$, where $V$ denotes the set of balls and $E$ the paths between them. Each ball $i\in V$ is assigned a reward $p_{i}$, and each edge $(i,j)\in E$ has an associated travel cost $c_{ij}$. Starting at a designated initial state $s\in V$, the goal is to determine a path $T$ through a subset of nodes $S\subseteq V$ that maximizes the cumulative rewards $\sum_{i\in S}p_{i}$, while ensuring that the total travel cost $\sum_{(i,j)\in T}c_{ij}$ does not exceed a predefined budget $C_{\text{max}}$. Formally, this optimization problem is expressed as:
@@
$\sum_{j\in V}x_{ij}-\sum_{j\in V}x_{ji}=\begin{cases}1&\text{if }i=s,\\ 0&\text{otherwise},\end{cases}\quad\forall i\in V,$ (5c)
$x_{ij}\in\{0,1\},\quad\forall(i,j)\in E,$ (5d)
$y_{i}\in\{0,1\},\quad\forall i\in V.$ (5e)
@@
A.5 Exemplary Environment Interaction
Targeted-Exploration Prompting Your past episodes: Episode 1: You see: dodger_blue door, tangerine door >dodger_blue door Reward: 0 You see: apricot door, dodger_blue door, cerulean door, honeydew door >cerulean door Reward: 0 You see: rosewood ball, teal door, cerulean door >rosewood ball Reward: 3 You see: teal door, cerulean door >teal door Reward: 0 You see: turquoise ball, magenta door, teal door, khaki ball >turquoise ball Reward: 2 You see: magenta door, teal door, khaki ball >khaki ball Reward: 3 You see: magenta door, teal door You are controlling an agent in an unknown world. Over a total of 20 episodes, you can interact with objects in the environment. You have 8 door interactions per episode but can pick up three balls, keys, or boxes. Based the current and past trials, explore the environment to collect information that may help to become better at maximizing the reward. Current episode: You see: dodger_blue door, tangerine door >tangerine door Reward: 0 You see: midnight_blue ball, tangerine door >midnight_blue ball Reward: 5 You see: tangerine door >tangerine door Reward: 0 You see: tangerine door, dodger_blue door Which object do you want to interact with next? Reply with one object enclosed with <and >,e.g. <door>. What is your next action?
Figure 8: Prompt Template
Appendix B Additional Results
B.1 Measuring Exploration with Large Language Model Exploitation
@@
(a) LLM Exploration Gap
(b) Optimal Exploration Gap (ours)
Figure 9: We find that LLM exploitation is similarly hard to exploration and cannot reliably measure exploration progress. Instead, we propose to rely on exact methods to fairly evaluate exploration.
B.2 Learning Curves
@@
Figure 10: Overall
@@
Figure 11: 4x4 Grid
@@
Figure 12: 5x5 Grid
@@
Figure 13: 7x7 Grid
B.3 Total Gap Decomposition
@@
Table 3: Total gap decomposition averaged over 4x4, 5x5, and 7x7 grids
@@
Table 4: Total gap decomposition of the 4x4 grids
@@
Table 5: Total gap decomposition of the 5x5 grids
@@
Table 6: Total return gap decomposition of the 7x7 grids
B.4 Other Statistics
@@
Table 7: Other statistics averaged over 4x4, 5x5, and 7x7 grids
@@
Table 8: Other statistics of the 4x4 grids
@@
Table 9: Other statistics of the 5x5 grids
@@
Table 10: Other statistics of the 7x7 grids
B.5 What Impact do Agent Instructions have on Exploration?
We demonstrate how our evaluation framework can be used to analyze the impact of prompt-engineering on exploration. All prompts used to evaluate the impact of prompt-engineering on exploration behaviors is illustrated in figure 14. To test the limits of LLM exploration, we design two prompts besides the task-oriented exploration (middle) from previous experiments. One of the prompts is written to analyze if agents can do full exploitation of their history (left), while the other should make agents explore aimlessly without mentioning a task.
Figure 14: Our prompts on a scale from optimal exploitation to exploration.
@@ B.5 What Impact do Agent Instructions have on Exploration?
Figure 15: Impact of the agent instruction on open-source model behavior.
(a) GPT-4o
(b) o1-mini
(c) Gemini 1.5 Pro
(d) Claude 3.5 Sonnet
Figure 16: Impact of the prompt on closed model behavior
B.5.1 Soft Exploration Lower Bound
@@
Table 11: Return gap decomposition of the 7x7 grids
@@
Table 12: Percentage change in the return gap decomposition of the 7x7 grids
@@
Table 13: Other statistics of the 7x7 grids
@@
Table 14: Percentage change in the other statistics of the 7x7 grids
B.5.2 Soft Exploration Upper Bound
@@
Table 15: Return gap decomposition of the 7x7 grids
@@
Table 16: Percentage change in the return gap decomposition of the 7x7 grids
@@
Table 17: Other statistics of the 7x7 grids
@@
Table 18: Percentage change in the other statistics of the 7x7 grids

### v2

@@ 6 Conclusion
In conclusion, we proposed a novel evaluation framework for LLM-based agents that isolates state-space exploration through a decomposition based on the optimal exploitation return. Our approach incorporates adaptations of popular environments from the RL domain, moving beyond multi-armed bandits. We argue that developing systems capable of both intelligent exploitation and exploration is a pivotal step toward creating truly generalist agents. During the presentation of our framework, we highlighted various research directions for which it offers a robust testbed.
Future work could extend our framework with more complex environments, featuring, for example, randomization and stochasticity. A limitation of our domain is that rewards are randomly spread and high return correlates with high state-space coverage. By developing environments that combine complex exploration with underlying rules, domain knowledge and different tasks, our framework could be used to test more intelligent exploration. Hereby, benchmarks such as WebArena Zhou et al. (2023) could serve as a starting point. Exploitation and exploration are both difficult problems. Therefore, it is essential to continue evaluating them in isolation.
Appendix A Evaluation Framework
@@
A.1 Dark Treasure Rooms
Figure 4: $4\times 4$ variations of Dark Treasure Rooms.
Figure 5: $5\times 5$ variations of Dark Treasure Rooms.
Figure 6: $7\times 7$ variations of Dark Treasure Rooms.
A.2 Maze
Figure 7: The maze layouts used in the experiments. Our variations are created using Kruskal’s algorithm. The starting position is the center. Rewards are sampled uniformly from $U[1,10]$ and distributed randomly across the maze. The agent can take a maximum of 15 steps in each episode.
A.3 Additional Environment Details
@@
Implementation. Our environments are built upon the minigrid framework. To address the limitations of the low-level action and observation spaces, we use an abstraction wrapper function. This function transforms raw observations and actions into higher-level abstractions, allowing agents to reason more effectively about interactions with objects and doors within a room. By abstracting low-level details, the evaluation focuses more on the agent’s exploration strategy. The abstraction wrapper defines the state as a discrete set of available objects and doors within the current room. Each room is uniquely configured to ensure that the agent can differentiate between them. The action space is simplified to interactions with specific objects or doors (e.g., choosing to open a door or collect an object). Once the agent selects an action at this abstract level, the wrapper translates it into a sequence of low-level actions using a predefined low-level policy to execute the selected behavior. Our wrapper can be seen as bridging the minigrid framework to a representation similar to text-environments like TextWorld, but minimizes narrative understanding and semantic associations. During the environment creation, we randomly drop doors and walls for Dark Treasure Rooms with probability $p_{\text{drop}}=0.01$ to increase the difficulty. We experimented with different model temperatures settings but did not find significant differences. Therefore, we set the hyperparameter, due to technical constraints, to $1.0$ for o1-mini and $0.1$ for the other models.
Source code. We open-source the code for our evaluation and the minigrid-wrapper after the peer-review. We also publish our framework on Huggingface Spaces to facilitate experimentation.
A.4 Prompt Template
In-Context Learning Agent Your past episodes: <HISTORY> You are controlling an agent in an unknown world. Over a total of 20 episodes, you can interact with objects in the environment. <INSTRUCTION> You have <MAX STEPS> door interactions per episode but can pick up three balls, keys, or boxes. Current episode: <EPISODIC HISTORY> Which object do you want to interact with next? Reply with one object enclosed with <and >, e.g. <door>. What is your next action?
Figure 8: The template used to prompt the LLMs. <History> is replaced with past episodes. <MAX STEPS> is set to 15 for the mazes and to the number of steps required to reach the furthest room in Dark Treasure Rooms. <EPISODIC HISTORY> shows the current agent-environment interactions. <INSTRUCTION> is where we command the agent to perform a specific task. In total we compare three different instructions. See appendix D for the specific details.
A.5 Exemplary Environment Interaction
Targeted-Exploration Prompting Your past episodes: Episode 1: You see: dodger_blue door, tangerine door >dodger_blue door Reward: 0 You walk through the door. You see: apricot door, dodger_blue door, cerulean door, honeydew door >cerulean door Reward: 0 You walk through the door. You see: rosewood ball, teal door, cerulean door >rosewood ball Reward: 3 You see: teal door, cerulean door >teal door Reward: 0 You walk through the door. You see: turquoise ball, magenta door, teal door, khaki ball >turquoise ball Reward: 2 You see: magenta door, teal door, khaki ball >khaki ball Reward: 3 You see: magenta door, teal door You are controlling an agent in an unknown world. Over a total of 20 episodes, you can interact with objects in the environment. You have 8 door interactions per episode but can pick up three balls, keys, or boxes. Based the current and past trials, explore the environment to collect information that may help to become better at maximizing the reward. Current episode: You see: dodger_blue door, tangerine door >tangerine door Reward: 0 You walk through the door. You see: midnight_blue ball, tangerine door >midnight_blue ball Reward: 5 You see: tangerine door >tangerine door Reward: 0 You walk through the door. You see: tangerine door, dodger_blue door Which object do you want to interact with next? Reply with one object enclosed with <and >,e.g. <door>. What is your next action?
Figure 9: An exemplary prompt after multiple agent-environment interactions. In the first episode, the agent agent traveled through multiple doors and collected three balls. In the current trial, the agent takes a different path and discovers another ball which leads to more reward than the others.
A.6 Optimal Exploitation
@@
We formulate the computation of the optimal exploitation return $R^{opt}$ as an instance of the Orienteering Problem. We represent the explored balls in the state-space of history $h_{t}$ as a graph $G=(V,E)$, where $V$ denotes the set of balls and $E$ the paths between them. Each ball $i\in V$ is assigned a reward $p_{i}$, and each edge $(i,j)\in E$ has an associated travel cost $c_{ij}$. Starting at a designated initial state $s\in V$, the goal is to determine a path $T$ through a subset of nodes $S\subseteq V$ that maximizes the cumulative rewards $\sum_{i\in S}p_{i}$, while ensuring that the total travel cost $\sum_{(i,j)\in T}c_{ij}$ does not exceed a predefined budget $C_{\text{max}}$ and a maximum of $C_{\text{balls}}$ balls is collected. Formally, this optimization problem is expressed as:
@@
$\sum_{i\in V}y_{i}\leq C_{\text{balls}},s$ (5c)
$\sum_{j\in V}x_{ij}-\sum_{j\in V}x_{ji}=\begin{cases}1&\text{if }i=s,\\ 0&\text{otherwise},\end{cases}\quad\forall i\in V,$ (5d)
$x_{ij}\in\{0,1\},\quad\forall(i,j)\in E,$ (5e)
$y_{i}\in\{0,1\},\quad\forall i\in V.$ (5f)
@@
A.7 Measuring Exploration with Large Language Model Exploitation
@@
(a) LLM Exploitation
(b) Optimal Exploitation
Figure 10: Comparison of LLM and our optimal exploitation. We find that LLM exploitation (left) is similarly hard to exploration and cannot reliably measure exploration progress. Instead, we propose to rely on exact methods (right) to fairly evaluate exploration.
Appendix B Results - Dark Treasure Rooms
B.1 Learning Curves
@@
Figure 11: Averaged over Dark Treasure Rooms $4\times 4$, $5\times 5$, and $7\times 7$.
@@
Figure 12: 4 $\times$ 4 Dark Treasure Rooms.
@@
Figure 13: 5 $\times$ 5 Dark Treasure Rooms.
@@
Figure 14: 7 $\times$ 7 Dark Treasure Rooms.
B.2 Total Return Gap Decomposition
Table 3: Normalized total gap decomposition averaged over $4\times 4$, $5\times 5$, and $7\times 7$ Dark Treasure Rooms. The numbers in parentheses show the fraction that exploitation and exploration gaps make up of the total return gap.
@@ B.2 Total Return Gap Decomposition
Table 4: Normalized total gap decomposition of $4\times 4$ Dark Treasure Rooms. The numbers in parentheses show the fraction that exploitation and exploration gaps make up of the total return gap.
@@
Table 5: Normalized total gap decomposition of $5\times 5$ Dark Treasure Rooms. The numbers in parentheses show the fraction that exploitation and exploration gaps make up of the total return gap.
@@
Table 6: Normalized total gap decomposition of $7\times 7$ Dark Treasure Rooms. The numbers in parentheses show the fraction that exploitation and exploration gaps make up of the total return gap.
@@
B.3 Other Statistics
Table 7: Other statistics averaged over $4\times 4$, $5\times 5$, and $7\times 7$ Dark Treasure Rooms.
@@ B.3 Other Statistics
Table 8: Other statistics of $4\times 4$ Dark Treasure Rooms.
@@
Table 9: Other statistics of the $5\times 5$ Dark Treasure Rooms.
@@
Table 10: Other statistics of $7\times 7$ Dark Treasure Rooms.
@@
Appendix C Results - Maze
C.1 Learning Curves
(a) Agent Return
(b) Exploration Gap
(c) Agent Return
(d) Exploration Gap
Figure 15: Learning curves averaged over all maze variations.
C.2 Total Return Gap Decomposition
Table 11: Normalized total gap decomposition averaged over all maze variations. The numbers in parentheses show the fraction that exploitation and exploration gaps make up of the total gap.
Last Episode Mean
<30B Mistral (7B) 1.00 0.30 (0.30) 0.70 (0.70) 0.97 0.26 (0.27) 0.71 (0.73)
Llama 3.1 (7B) 0.99 0.28 (0.29) 0.70 (0.71) 0.96 0.17 (0.18) 0.79 (0.82)
Gemma 2 (9B) 0.99 0.17 (0.17) 0.81 (0.83) 0.92 0.11 (0.11) 0.81 (0.89)
Mistral Small (22B) 0.95 0.29 (0.31) 0.66 (0.69) 0.95 0.16 (0.17) 0.80 (0.83)
Gemma 2 (27B) 0.52 0.06 (0.09) 0.46 (0.91) 0.54 0.05 (0.08) 0.49 (0.92)
>30B Llama 3.3 (70B) 0.80 0.28 (0.29) 0.53 (0.71) 0.86 0.30 (0.31) 0.56 (0.69)
Mistral Large (123B) 0.48 0.21 (0.41) 0.27 (0.59) 0.51 0.17 (0.32) 0.34 (0.68)
Closed GPT-4o 0.61 0.22 (0.42) 0.39 (0.58) 0.64 0.18 (0.32) 0.46 (0.68)
GPT-o1-mini 0.68 0.44 (0.63) 0.24 (0.37) 0.70 0.36 (0.52) 0.34 (0.48)
Gemini 1.5 Pro 0.83 0.39 (0.54) 0.44 (0.46) 0.74 0.24 (0.35) 0.50 (0.65)
Claude 3.5 Sonnet 0.50 0.06 (0.10) 0.45 (0.90) 0.56 0.06 (0.10) 0.50 (0.90)
Baseline Random Walk 0.96 0.51 (0.53) 0.45 (0.47) 0.95 0.29 (0.31) 0.65 (0.69)
C.3 Other Statistics
Table 12: Other statistics averaged over all maze variations.
<30B Mistral (7B) 7.13 0.00 12.17 0.97 9.81
Llama 3.1 (7B) 6.83 0.33 19.78 0.94 26.11
Gemma 2 (9B) 4.33 0.27 10.75 0.97 5.00
Mistral Small (22B) 8.03 1.17 18.24 0.93 39.00
Gemma 2 (27B) 12.17 10.70 25.06 0.93 10.17
>30B Llama 3.3 (70B) 10.90 4.40 34.27 0.91 17.00
Mistral Large (123B) 16.73 11.97 48.09 0.87 21.04
Closed GPT-4o 13.93 9.03 38.24 0.90 23.85
GPT-o1-mini 17.40 7.33 57.34 0.85 31.67
Gemini 1.5 Pro 12.90 3.97 28.76 0.90 25.36
Claude 3.5 Sonnet 12.60 11.30 35.09 0.91 17.31
Baseline Random Walk 12.60 0.83 42.77 0.88 54.67
Appendix D What Impact do Agent Instructions have on Exploration?
Our experiments aim to demonstrate how our evaluation framework can be used to analyze the impact of prompt-engineering on the exploration capability. In this context, we designed two agent instructions to test the limits of our agents. We compare their total return decomposition to our task-oriented exploration. All prompts used to evaluate the impact on the exploration are illustrated in Figure 16.
Soft Lower Bound: In order to create a lower exploration bound, we instruct the agent to strictly stick to known rewards. This way, we investigate if agents can perform constant exploitation. The prompt is shown on the left of Figure 16. We expect a continuously low exploitation and a high exploration gap.
Soft Upper Bound: We designed another prompt to measure the upper exploration performance. The instruction is supposed to direct the agent towards an aimless exploration as we do not mention a task. This also allows us to investigate if goal orientation hinders exploration. As high optimal exploitation returns correlate with state-space coverage in our environments, we expect this setting to achieve the lowest exploration gap.
Figure 16: The agent instructions used for the task-oriented exploration as well as our soft lower and upper bounds. The Figure displays the prompts on a hypothetical and subjective continuum which ranges from strict exploitation to exploration.
(a) GPT-4o
(b) o1-mini
(c) Gemini 1.5 Pro
(d) Claude 3.5 Sonnet
Figure 17: Impact of the agent instruction on the total return gap decomposition for closed models.
@@ Appendix D What Impact do Agent Instructions have on Exploration?
Figure 18: Impact of the agent instruction on the total return gap decomposition for closed models.
D.1 Soft Exploration Lower Bound
Table 13: Normalized total return gap decomposition of $7\times 7$ Dark Treasure Rooms. The numbers in parentheses show the fraction that exploitation and exploration gaps make up of the total return gap.
@@ D.1 Soft Exploration Lower Bound
Table 14: Percentage change in the total return gap decomposition of $7\times 7$ Dark Treasure Rooms.
@@
Table 15: Other statistics of $7\times 7$ Dark Treasure Rooms.
@@
Table 16: Percentage change in the total return gap decomposition of $7\times 7$ Dark Treasure Rooms.
@@
D.2 Soft Exploration Upper Bound
Table 17: Normalized total return gap decomposition of $7\times 7$ Dark Treasure Rooms. The numbers in parentheses show the fraction that exploitation and exploration gaps make up of the total return gap.
@@ D.2 Soft Exploration Upper Bound
Table 18: Percentage change in the total return gap decomposition of $7\times 7$ Dark Treasure Rooms.
@@
Table 19: Other statistics of $7\times 7$ Dark Treasure Rooms.
@@
Table 20: Percentage change in the total return gap decomposition of $7\times 7$ Dark Treasure Rooms.
@@
Appendix E Additional Future Implications
Besides the research directions in the main paper, we believe the following paths are interesting to investigate in combination with our evaluation framework:
Memory architectures.    Many agent systems currently focus on developing memory architectures Wang et al. (2024); Anokhin et al. (2024) that load the LLM context dynamically. Testing such architectures for systematic exploration with an optimal exploitation, should be an important dimension in the evaluation. Further, building a meta memory that allows to reuse exploration-specific information across tasks and environments may be a fruitful path.
Improving sequence understanding. Similarly to Paglieri et al. (2024), we found that agents increasingly produce invalid actions and misunderstand longer sequences of past actions. Developing agents specifically made for long-sequence understanding is crucial for future LLM-based agents. It would be interesting to see if LLMs finetuned on sequential data Kim et al. (2024); Zeng et al. (2023); Song et al. (2024) also exhibit better exploration capabilities.